a stone thrown from the top of the building is given initial velocity off 20m/s streight upwards, the stone is launched 50m above the ground and the stone just misses the edge of the roof on its way down using ta = 0

1
Expert's answer
2022-06-07T17:34:21-0400

"t_1=\frac vg=2~s,"

"s=\frac{v^2}{2g}=20~m,"

"t_2=\sqrt\frac{2(s+h)}g=3.7~s,"

"t=t_1+t_2=5.7~s."